OpenStudy (anonymous):

PLEASE HELP!! write in equation in slop-intercept form for the line that passes through (4,6)and is perpendicular to line given by y=2x+3

OpenStudy (anonymous):

general linear equation y=mx+c m=gradient c=y-intercept perpendicular tells us that we have to use this formula m1 x m2 = -1 m1=gradient of equation given m1=2 m2=gradient of unknown equation we want to find 2 x m2 = -1 m2=-1/2 now substitute (4,6) and gradient m=-1/2 into the general equation y=mx+c to find the c value and then re-write the whole equation.